40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 6960     Entry time: Wed Jul 11 13:36:58 2012
Author: yuta 
Type: Update 
Category: SUS 
Subject: OSEM and oplev spectra of optics 

Below is angular spectra of every suspended core optics.
As you can see, there's a peak at 3.3 Hz for BS and PRM angular motion. Compared with other optics, they look large.

I briefly checked suspension filters and found that BounceRoll filters and f2a filters are not turned on for BS.
I checked elog and there was no reason for them to be off, so I turned them on. It didn't change angular spectra very much, though.

I'm going to check BS suspension damping and see where 3.3 Hz peak comes from.

Note that oplev quadrant sums are different for every optics, so we can't simply compare angular motion between optics from OLPIT/OLYAW. But for OSEMs, there are "cnt2um" which calibrate sensor outputs into um. and input matrix should be normalized, so we can compare SUSPIT/SUSYAW with other optics.

I centered all oplevs to do this measurement.
Quadrant sum (C1:SUS-XXX_OLSUM) for each optic now is

ITMX   ITMY   ETMX   ETMY     BS    PRM    SRM
2456  14630   1476  14885   3650   4302   2937   (counts)


OLPITYAW.png   SUSPITYAW.png

ELOG V3.1.3-